OpenStudy (cometailcane):

If a 1.85 g mass of zinc produces 475 mL of gas and your balloon weighs 0.574 g and the room temperature is 20.5°C. Calculate the amount of zinc needed to produce enough gas to get your balloon airborne by adding 1 mL to the required balloon volume so that its density is less than that of the surrounding air.

OpenStudy (cometailcane):

I came up with a required balloon volume of 476 before adding the extra 1 mL.

OpenStudy (aaronq):

How did you come up with that? have you also attempted to obtain a value for the mass of Zn?

OpenStudy (cometailcane):

By using y=-0.0041x+1.289 which is the equation

OpenStudy (cometailcane):

and then V= mass/density V=0.574/1.20495

OpenStudy (cuanchi):

I think you still having problems with the units air density = 1.20495 kg/m^3 = 1.20495 x 10^3 g /m^3 V=0.574g/(1.20495 x 10^3 g/m^3) V= 0.477 m^3 = 477 L 477000 mL+ 1 mL = 477001 mL grams of Zn= 477001 mL x (1.85 g Zn / 475 mL) = ???? g of Zn